# Reinforcement Learning Finetunes Small Subnetworks in Large Language Models

### v1

@@ 5 Consistency of Subnetworks Across Seeds, Data, and Algorithms
We evaluate three settings: (1) varying the random seed alone, (2) varying the training data alone, and (3) changing the seed, data, and the RL algorithm altogether as a stress test. We conduct controlled experiments and all factors not under investigation are the same. Unless otherwise mentioned, all ablations were done with a batch size of 32, trained for one epoch with the base model Tulu-3-8B-SFT. When varying the training data, we switch between the Tulu preference dataset333allenai/llama-3.1-tulu-3-8b-preference-mixture and the PRIME rollout dataset444 PRIME-RL/Eurus-2-Rollout. It has model generations for math datasets alongside a label for correctness. To adapt the rollout dataset to DPO format, we select only positive samples, and pair it with a randomly sampled negative one. When varying the RL algorithm, we train a DPO model initialized from PRIME-RL/Eurus-2-7B-SFT and compare it to the PRIME-RL/Eurus-2-7B-PRIME model.
@@ 6 Why Do the Subnetworks Emerge?
    What factors contribute to the update sparsity observed in RL finetuning?
@@
Performing SFT before RL. A common design choice is to perform SFT on the same data as the subsequent RL Ouyang et al. (2022). However, as shown in Table 1, our findings extend to models such as DeepSeek-R1-Zero, which forgoes SFT entirely yet still exhibit high update sparsity. This suggests that SFT is not a main contributing factor to the update sparsity in RL finetuning.
Training duration. It is intuitive that with more gradient steps, a model is expected to drift more from the base model. Figure 5 shows how update sparsity changes during the training of PRIME. As training progresses, update sparsity gradually decreases but eventually shows signs of convergence to around 80%. Notably, DeepSeek-R1-Zero undergoes 8K training steps (numbers from Figure 2 in DeepSeek-AI et al. (2025)) using GRPO, over 20$\times$ more than PRIME, but shows a comparable update sparsity (86%). Therefore, we conjecture that training duration’s impact on update sparsity is more prominent during early training but gradually decreases as training progresses.
@@
While it is possible, though less likely, that all models we study, including DeepSeek-R1-Zero (86.0% update sparsity after 8K steps), are severely undertrained, and that the observed update sparsity would diminish with substantially more training. Nonetheless, we question the practicality of this hypothesis since it runs counter to the RL literature arguing against overtraining to prevent overfitting and improve generalization Fu et al. (2019).
Training on in-distribution data. Intuitively, when gradients are computed on sequences that the policy already assigns high probabilities to, little update to the parameters would be needed. We evaluated two scenarios: (1) rejection sampling, and (2) DPO on out-of-distribution data by not performing SFT pridor to RL. Since we’ve already learned from §3 that DPO with in-distribution data induces sparse updates while SFT (with out-of-distribution data) induces dense ones, this additional experiment can serve as a control group to isolate the factor of training on in- vs. out-of-distribution data.
As shown in Table 5, Our experiments reveal that SFT on in-distribution data produces sparse updates, while DPO with out-of-distribution data produces dense ones. Specifically, performing SFT with Qwen/Qwen2.5-Math-7B on rejection sampled in-distribution data yields around 90.0% update sparsity. This is reinforced by the examination of a previous work: RAFT++ (Xiong et al., 2024), which performs supervised finetuning with iterative rejection sampling, yields an update sparsity of 69.4%. In contrast, DPO on out-of-distribution data produces dense updates in zephyr-7b-beta models, with a 6.8% update sparsity. These findings suggest that training on in-distribution data is a major driver of update sparsity in not only RL, but also SFT.
Takeaway 4 Training on in-distribution data is a primary reason of update sparsity; KL-divergence regularization and gradient clipping have limited impact.
Table 5: RFT indicates rejection-sampling fine-tuning (Touvron et al., 2023; Dong et al., 2023), and RAFT++ is iterative RFT A comparative analysis across SFT and DPO as well as in- vs out-of-distribution training shows that in-distribution consistently produces sparse updates.
Model Method Sparsity (%) SFT/RL In-Dist
@@ 7 Limitations and Future Work
Because RL is computationally demanding, we choose to vary one factor at a time; yet the observed sparsity may actually result from complex interactions among many, an avenue future work should examine. Further, fully controlled experiments are computationally prohibitive, and we thus sometimes resort to resort to public checkpoints. While our experiments focus on language models, it would be interesting to explore the same questions for multimodal and diffusion models. Subsequent research could investigate methods for early identification of the sparse subnetwork and ways to leverage its structure for more efficient learning. Finally, our empirical findings invite a deeper theoretical analysis, with the goal of uncovering theoretical explanations for the update sparsity in RL.
@@ 8 Conclusion
Our study reveals that RL finetuning in LLMs updates only a sparse subnetwork constituting approximately 5%-30% of total parameters, leaving the rest unchanged. This sparsity emerges intrinsically, without explicit spasity promoting techniques such as regularization or structural constraints. Crucially, finetuning just this subnetwork in isolation reproduces the full model’s performance, aligning closely with original parameter values. For a given base model across different seeds, datasets and learning algorithms, a non-trivial portion of the subnetwork remains the same. Our findings highlight that learning from in-distribution samples while training is a key driver of this phenomenon, pointing towards more efficient and effective training strategies in RL-based finetuning of LLMs.

### v2

@@ 5 Consistency of Subnetworks Across Seeds, Data, and Algorithms
We evaluate three settings: (1) varying the random seed alone, (2) varying the training data alone, and (3) changing the seed, data, and the RL algorithm altogether as a stress test. We conduct controlled experiments and all factors not under investigation are the same. Unless otherwise mentioned, all ablations were done with a batch size of 32, trained for one epoch with the base model Tulu-3-8B-SFT. When varying the training data, we switch between the Tulu preference dataset333allenai/llama-3.1-tulu-3-8b-preference-mixture and the PRIME rollout dataset [^4] PRIME-RL/Eurus-2-Rollout. It has model generations for math datasets alongside a label for correctness. To adapt the rollout dataset to DPO format, we select only positive samples, and pair it with a randomly sampled negative one. When varying the RL algorithm, we train a DPO model initialized from PRIME-RL/Eurus-2-7B-SFT and compare it to the PRIME-RL/Eurus-2-7B-PRIME model.
@@ 6 Why Do the Subnetworks Emerge?
    What factors contribute to the update sparsity observed in RL finetuning? [^5] We conjecture that, for the sake of illustration, if one were to perform backpropagation manually on paper with unlimited numerical precision, the resulting parameter updates would be dense. In practice, however, modern computers rely on floating-point arithmetic with limited precision. As a result, updates with very small magnitudes (e.g., absolute values below $10^{-40}$) cannot be represented and are effectively discarded, hence the empirically observed sparsity of the parameter updates. Importantly, such near-zero updates appear to have negligible impact on model performance, as strong results have been achieved without them. The equivalent question we address in this section is: What factors contribute to these near-zero parameter updates during RL fine-tuning?
@@
Performing SFT before RL. A common design choice is to perform SFT on the same data as the subsequent RL ouyang2022traininglanguagemodelsfollow. However, as shown in Table 1, our findings extend to models such as DeepSeek-R1-Zero, which forgoes SFT entirely yet still exhibit high update sparsity. This suggests that SFT is not a main contributing factor to the update sparsity in RL finetuning.
[…]
@@
While it is possible, though less likely, that all models we study, including DeepSeek-R1-Zero (86.0% update sparsity after 8K steps), are severely undertrained, and that the observed update sparsity would diminish with substantially more training. Nonetheless, we question the practicality of this hypothesis since it runs counter to the RL literature arguing against overtraining to prevent overfitting and improve generalization pmlr-v97-fu19a.
Training on in-distribution data. Intuitively, when gradients are computed on sequences that the policy already assigns high probabilities to, little update to the parameters would be needed. We evaluated two scenarios: (1) rejection sampling, and (2) DPO on out-of-distribution data by not performing SFT prior to RL. Since we’ve already learned from §3 that DPO with in-distribution data induces sparse updates while SFT (with out-of-distribution data) induces dense ones, this additional experiment can serve as a control group to isolate the factor of training on in- vs. out-of-distribution data.
As shown in Table 5, Our experiments reveal that SFT on in-distribution data produces sparse updates, while DPO with out-of-distribution data produces dense ones. Specifically, performing SFT with Qwen/Qwen2.5-Math-7B on rejection sampled in-distribution data yields around 90.0% update sparsity. This is reinforced by the examination of a previous work: RAFT++ (xiong2024iterative), which performs supervised finetuning with iterative rejection sampling, yields an update sparsity of 69.4%. In contrast, DPO on out-of-distribution data produces dense updates in zephyr-7b-beta models, with a 6.8% update sparsity. These findings suggest that training on in-distribution data could be a major driver of update sparsity in not only RL, but also SFT.
Takeaway 4 We conjecture that training on in-distribution data could be a reason of update sparsity; KL-divergence regularization and gradient clipping have limited impact.
Table 5: RFT indicates rejection‐sampling fine‑tuning (touvron2023llama2openfoundation; dong2023raftrewardrankedfinetuning), and RAFT++ is iterative RFT A comparative analysis across SFT and DPO as well as in- vs out-of-distribution training shows that in-distribution consistently produces sparse updates.
Model Method Sparsity (%) SFT/RL In‐Dist
@@ 7 Limitations and Future Work
Because RL is computationally demanding, we choose to vary one factor at a time; yet the observed sparsity may actually result from complex interactions among many, an avenue future work should examine. Further, fully controlled experiments are computationally prohibitive, and we thus sometimes resort to resort to public checkpoints. While our experiments focus on language models, it would be interesting to explore the same questions for multimodal and diffusion models. Subsequent research could investigate methods for early identification of the sparse subnetwork and ways to leverage its structure for more efficient learning. Finally, our empirical findings invite a deeper theoretical analysis, with the goal of uncovering theoretical explanations for the update sparsity in RL. Lastly, while our observations generally hold, there are confounders (for eg. liu2025prorlprolongedreinforcementlearning). However the confounding reason is not trivial and is worth exploring as part of future work. Further to the best of our observations this phenomenon is largely applicable to all RL settings.
@@ 8 Conclusion
Our study reveals that RL finetuning in LLMs updates only a sparse subnetwork constituting approximately 5%-30% of total parameters, leaving the rest unchanged. This sparsity emerges without explicit spasity promoting techniques such as regularization or structural constraints. Crucially, finetuning just this subnetwork in isolation reproduces the full model’s performance, aligning closely with original parameter values. For a given base model across different seeds, datasets and learning algorithms, a non-trivial portion of the subnetwork remains the same. Our findings highlight that learning from in-distribution samples while training is a key driver of this phenomenon, pointing towards more efficient and effective training strategies in RL-based finetuning of LLMs.
9 Acknowledgments
We would like to thank Pavan Jayasinha, Cheng Wang, Abdul Waheed, Shivanshu Shekhar, Alexander Schwing, Gabriel Stanovsky, Roy Schwartz and Gokhan Tur for their valuable feedbacks and suggestions. Further we extend our gratitude to the members of ConvAI and Alta lab groups at UIUC for their thoughtful remarks on our draft. This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Sections 3-6 provides empirical evidence for the claims made in intro and abstract
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Separate dedicated limiation section included
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: The paper doesnt discuss theoretical proofs, most results are empirical
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: All relevant datasets, models are listed, all hyperparameters discussed in detail
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification:
  21. Guidelines:
  22. 6.
    Experimental setting/details
  23. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  24. Answer: [Yes]
  25. Justification: All details shared for reproducibility
  26. Guidelines:
  27. 7.
    Experiment statistical significance
  28. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  29. Answer: [No]
  30. Justification: Most accuracies are reported on deterministic setups
  31. 8.
    Experiments compute resources
  32. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  33. Answer: [Yes]
[…]
  35. 9.
    Code of ethics
  36. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  37. Answer: [Yes]
  38. Justification: Most experiments performed are with publicly released models, datasets and codebases.
  39. 10.
    Broader impacts
  40. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  41. Answer: [N/A]
  42. Justification: No immediate societal impacts of the work
  43. 11.
    Safeguards
  44. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  45. Answer: [N/A]
  46. Justification: No such risks
  47. 12.
    Licenses for existing assets
  48. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  49. Answer: [Yes]
  50. Justification: All works are cited.
  51. 13.
    New assets
  52. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  53. Answer: [N/A]
  54. Justification: Doesnt release new assets
  55. 14.
    Crowdsourcing and research with human subjects
  56. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  57. Answer: [N/A]
  58. Justification: Doesn’t involve crowdsourcing.
  59. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  60. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  61. Answer: [N/A]
  62. Justification: does not involve crowdsourcing nor research with human subjects
  63. 16.
    Declaration of LLM usage
  64. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  65. Answer: [N/A]
  66. Justification: Does not involve LLMs usage